# Pretraining in Actor-Critic Reinforcement Learning for Robot Motion Control

### v1

@@ 4.4 Pretraining the Proprioceptive Inverse Dynamics Model
We parameterize the PIDM with an MLP-based modular architecture, as shown in Figure 3. The model takes as input a history of actions $a_{t-K:t-1}$ and proprioceptive observations $x_{t-K:t+1}$ of length $K$. Both are passed through a dual-layer MLP encoder before being fed into the PIDM backbone, which is a 4-layer MLP. During pretraining, we give the model a desired delta-state $\Delta x_{t+1}^{*}$ to achieve in the next time step. We then use an $L1$ loss to supervise the PIDM to output the required action $a_{t}$ to reach the target future state $x_{t+1}^{*}$. The pretraining dataset is also augmented with symmetry transformations, as defined by Mittal et al. (2024) or Byun & Perrault (2024), and observation noise to improve robustness and increase sample diversity (see Appendix A.4). It is important to note that the PIDM model does not have access to privileged information. The action and proprioception histories provide indirect observability of contact states, the domain randomization variables during training (e.g., mass and friction randomization), and of random forces being applied to the robot. This knowledge is crucial to obtaining an accurate prediction of the system’s dynamics.
@@ 4.5 Warm-starting Reinforcement Learning
Integrating PIDM into actor-critic networks: The pretrained PIDM is integrated into both the actor and critic networks. As shown in Figure 3, we first remove the Delta encoder and substitute it with a randomly initialized Intention encoder that processes the complete task-specific observation. The Intention encoder now only needs to learn an embedding-based representation of the task-specific delta target state $\Delta x_{t+1}^{*}$, which can be preprocessed by the pretrained PIDM backbone. Meanwhile, the original output-layer (Prediction decoder) is removed, and the concatenated outputs of the PIDM backbone and Intention encoder are passed in to a randomly initialized Action synthesizer that synthesizes the final action $a_{t}$.
@@
Intact RL setup: Except for the architectures of the actor and critic networks and the way the weights are initialized, our method does not require any modifications to either the POMDP (reward, curriculum design, observations, actions, and terminations) or to the PPO hyperparameters. The task-dependent Intention encoder and Action synthesizer can adapt to any configuration and dimension of the input and output. Therefore, the feasibility of handling arbitrary tasks is not limited. Every parameter in the pretrained PIDM remains trainable during the RL process. In this way, we allow task-specific dynamics to be learned during policy optimization, which eases the burden of attempting to exhaustively cover all possible transitions in the pretraining dataset.
@@ 5.1 Reinforcement Learning Tasks
We test our method on seven distinct tasks with the same ANYmal-D (Hutter et al., 2016) quadruped embodiment within Isaac Lab (Mittal et al., 2023):
  • Two blind tasks without access to any exteroception: velocity-commanded locomotion (Rudin et al., 2022) and pedipulation (Arm et al., 2024).
  • Five perceptive tasks with access to height scans from the parkour suite (Hoeller et al., 2023): parkour walk, climb up, climb down, crouch, and jump.
Results are averaged over five runs with different random seeds, except for some individual cases mentioned in Appendix Table 8. Despite the diverse philosophies in reward and curriculum design across these tasks, the network architectures used are similar in that the actor and critic networks are both 4-layer MLPs.
The compactness of the architectures can be attributed to the fact that the trained policy networks are expected to be deployed on real mobile hardware and be reactive at a high frequency (typically $50\sim 200$ Hz). Our proposed architecture has approximately $4\times$ the number of parameters due to the inclusion of state history and the need to cover a larger initial state space in pretraining, compared to task-specific policies that can immediately hyper-specialize.
@@ 5.2 Dynamics Knowledge Probing in Vanilla RL Policy Networks
(b) Final prediction errors after convergence.
Figure 4: Experiment to probe dynamics knowledge in vanilla policy networks. We analyze the correlation between internal layer states and the future joint state $q_{t+1}$, based on the predicted action $a_{t}$. The zero-order extrapolation in (b) is a reference of the accuracy of always predicting $q_{t}=q_{t+1}$.
As a means of empirically studying our initial Hypothesis 1, we examine five different checkpoints of the pedipulation (Arm et al., 2024) policy network from the RL process, taken at 1000 iteration intervals. For each checkpoint, we collect observation-action pairs $(o_{t},a_{t})$ from policy rollouts during training. We then execute the mean predicted action in simulation and observe the resulting change in joint angles $\Delta q_{t+1}$. Additionally, we collect the corresponding intermediate representations of all three hidden layers, as illustrated in Figure 4(a).
We investigate the correlation between the network’s internal representation and $\Delta q_{t+1}$, by fitting a lightweight MLP per layer to regress the consequences of the predicted action. The final prediction error, shown in Figure 4(b), is then inversely proportional to how well the network understands at a specific layer what the consequences of its action will be. On the horizontal axis, we see that during training, the model gains an understanding of dynamics. In the vertical direction, the correlation between the future state diminishes as we progress deeper into the network. This highlights our analogy to classic control from Section 4.1, where the model first forms an intent on the target state and subsequently computes the inverse dynamics to determine the required action. For more details on the experiment, we refer the readers to Appendix A.2.
@@ 5.3 Pretraining the Proprioceptive Inverse Dynamics Model
Figure 5: PIDM training and dataset analysis: For the pretraining dataset we visualize its (a) coverage (green) compared to the initial exploration stages in RL (red and blue) using an UMAP projection and (b) the sample distribution of absolute action magnitudes $|a_{t}-q_{t}|$ over different terrains. Finally, in (c) we show the resulting PIDM accuracy across the entire action range as absolute joint errors $|\hat{q}_{t+1}-q_{t+1}|$ and also normalized by the action magnitude.
We first analyze the quality of the data collected using the exploration-based strategy outlined in Section 4.3. In addition to the previously mentioned standard data augmentations (e.g., mass randomization, random noise, symmetry), we collect data on both flat and basic rough terrain generated with Perlin noise (Miki et al., 2022; Lee et al., 2020). However, in Figure 5(a) we only plot samples from the flat-terrain environments to maintain consistency with the pedipulation and locomotion tasks, which are trained only on flat terrain. Using UMAP (McInnes et al., 2018), we project the proprioceptive observations $x$ for our collected dataset into 2D, and the observations from the first $100$ iterations of RL training for pedipulation and locomotion. We can thus validate that we obtain good coverage of, and beyond, the initial stages of the RL training process. This aligns with the goals outlined in Section 4.2.
The PIDM is pretrained as described in Section 4.4. We use a total of $5.3$ million samples for training and a similarly sized, disjoint validation set. For plotting purposes only, we randomly select $1,000$ validation samples and consider only the $12$ joint angles $q$ of the quadruped robot. Figure 5(b) shows the distribution of the action magnitude, i.e. the magnitude in radians of the commanded change in angle. Figure 5(c) shows the final prediction accuracy of the trained PIDM. We show both the absolute error and the normalized error, which is the error expressed as a fraction of the action magnitude. We achieve a normalized error of around $40\%\sim 50\%$, with a minimum error of $\sim 0.1$ radians for small actions. This indicates the considerable difficulty in training a PIDM with high accuracy, which we attribute to the vast transition space, partial observability, and the lack of inductive bias in MLPs. For a detailed discussion, see Appendix A.4. Although this may not seem sufficiently accurate for an inverse-dynamics model, we will demonstrate in the following subsection that a pretrained module of such accuracy can already significantly enhance RL training.
@@ 5.4 Quantitative experiments
Metric Method Locomotion Pedipulation Parkour Climb Climb Crouch Jump Avg.
Walk Up Down
Final perf. (%, $\uparrow$) Vanilla MLP +0.5 +0.2 -0.8 0.0 +6.5 +1.8 +11.1 +2.8
increase PIDM (Pretrained) +10.1 +6.3 +0.7 0.0 +27.7 +1.8 +5.9 +7.5
Num. iters. (%, $\downarrow$) Vanilla MLP -28.7 +5.0 -18.7 +22.5 -11.4 -29.5 -53.0 -16.3
to converge PIDM (Pretrained) -33.1 -42.0 -35.3 -20.6 -43.2 -57.3 -49.3 -40.1
@@
Figure 6: Evolution of the main performance metric during training for four of the tasks we use to evaluate. The shaded area represents the variance across five differently seeded trainings.
In each experiment, we compare three methods: (i) the vanilla 4-layer MLP, (ii) our PIDM architecture with randomly initialized weights, and (iii) our PIDM architecture with pretrained weights. The utility of our method (i.e. using pretrained weights) is indicated by the comparison between (ii) and (iii). The performance of the vanilla 4-layer MLP is included only as a reference. To note is that we did not tune the learning parameters (learning rate, entropy coefficient, etc.) of the tasks, which were chosen for optimal performance of the original vanilla MLP. We merely used our architecture as a drop-in replacement. Therefore, we suspect that the performance of our method can be further improved with additional tuning of these hyperparameters, which would have happened if the problem design had used our proposed architecture as a starting point.
@@
The selection of the main performance indicator varies across tasks (for details, see Appendix A.6). The results across all seven tasks are presented in Table 1. For some tasks, we also plot the evolution of the main performance indicator during training in Figure 6. The PIDM architecture with random weight initialization, PIDM (Random Init), generally lags behind the vanilla MLPs due to a larger model size and input dimension (inclusion of history). However, with the proposed pretraining strategy, PIDM (Pretrained) not only consistently outperforms PIDM (Random Init) in all metrics, but also significantly surpasses the performance of the vanilla MLP in $6/7$ tasks. Matching the performance of the MLP with our PIDM architecture is a secondary goal that we believe could be achieved by exhaustive tuning of the model and proprioceptive input. The key takeaway is the comparison between the randomly initialized and pretrained architectures. Compared with the vanilla MLP, PIDM (Pretrained) demonstrates an average improvement of 4.74.7% on final performance, and 23.823.8% on sample efficiency. When compared with PIDM (Random Init), the proposed PIDM (Pretrained) showcases an improvement of 7.57.5% on final performance, and enhances sample efficiency by a margin of 40.140.1%. We also note that despite the PIDM never having experienced the complex terrains used in the parkour task (see Figure 1), it quickly adapts to the new task-specific dynamics during RL training.
@@ 5.5 Ablations
Metric Method Climb Climb
@@
Metric Method Climb Climb
@@
We also perform ablations to motivate some of our design choices. First, in Table 2 we analyze the initialization strategy of the actor and critic networks. We note that only pretraining either the actor or critic generally still improves mean performance, but at the cost of increased instability, underscored by larger variation across runs (see Figure 10 in the Appendix). Second, we also ablate the data source used to pretrain the PIDM. We compare using exploration-based data versus samples from the initial stages of RL training a policy (in this case, pedipulation, as shown in Table 3. Notably, both datasets significantly outperform random initialization. Our approach provides an extra margin of improvement and theoretically adapts better to downstream tasks without overcomplicating the pipeline, as pretraining is performed only once.
@@ 5.6 Weight Update Magnitude
Figure 7: Network weight update magnitude comparison in the PIDM structured actor network of the pedipulation during RL training. In each submodule, the update of each linear layer weight is indicated by the mean absolute change per parameter, which is then averaged over all layers.
We compare the network weight update magnitudes between PIDM (Pretrained) and PIDM (Random Init) during the first 100 iterations of RL in Figure 7. We find that not only does the model exhibit smaller updates per iteration in the pretrained PIDM backbone, but this also results in smaller updates in the randomly initialized Intention encoder and Action synthesizer. This finding suggests that our pretrained weights lie closer to the desired local minimum and is an indicator that the optimization process can properly leverage this fact. For more examples, see Figure 11 in Appendix.
@@ 6 Conclusion
To summarize, we have presented a method for warm-starting the RL training process in actor-critic algorithms, targeted for robotic motion control. Our proposed approach leverages a network architecture based on a Proprioceptive Inverse Dynamics Model (PIDM) that is pretrained using exploration-based data from a specific robot embodiment. In contrast to related work, our modeling of the problem addresses the sim-to-real gap in robotics by working with noise and other domain randomization techniques. Our modular architecture functions as a drop-in replacement, without hyperparameter tuning, for any task on the pretrained robot embodiment. We demonstrate on seven diverse RL tasks with a quadrupedal robot that we can improve the final performance by $\mathbf{7.5}$%, and enhance sample efficiency by $\mathbf{40.1}$%. We also provide ablation studies to motivate our design choices and numerous empirical insights into the inner workings of our method. Future work can focus on optimizing model design to reduce the network size further and incorporating network architectures that are more adept at working with time-series data.

### v2

@@ 4.4 Pretraining the Proprioceptive Inverse Dynamics Model
We parameterize the PIDM with an MLP-based modular architecture, as shown in Figure 3. The model takes as input a history of actions $a_{t-K:t-1}$ and proprioceptive observations $x_{t-K:t+1}$ of length $K$. Both are passed through a dual-layer MLP encoder before being fed into the PIDM backbone, which is a 4-layer MLP. During pretraining, we give the model a desired delta-state $\Delta x_{t+1}^{*}$ to achieve in the next time step. We then use an $L1$ loss to supervise the PIDM to output the required action $a_{t}$ to reach the target future state $x_{t+1}^{*}$. The pretraining dataset is also augmented with symmetry transformations, as defined by mittal2024symmetry or byun2024symmetric, and observation noise to improve robustness and increase sample diversity (see Appendix A.4).
The necessity of including a history of proprioception for PIDM is mainly due to the absense of terrain information and contact state in proprioception, and the presence of noise and domain randomization techniques in the training process (during both exploration-based data collection and training of task-specific policies). Therefore, it would be inappropriate to fit the PIDM with only one single frame of current proprioceptive state, due to the fact that one certain proprioception can be observed in a range of actual full states in the POMDP. The action and proprioception histories can provide indirect observability of contact states, the domain randomization variables during training (e.g., mass and friction randomization), and of random forces being applied to the robot (ji2022concurrent; portela2025whole). This knowledge is crucial for mastering the system’s dynamics. Meanwhile it is important to note that the PIDM model does not have access to privileged information.
@@ 4.5 Warm-starting Reinforcement Learning
Integrating PIDM into actor-critic networks: The pretrained PIDM is integrated into both the actor and critic networks. As shown in Figure 3, for the actor, we first remove the Delta Encoder and substitute it with a randomly initialized Intention Encoder that processes the complete task-specific observation. The Intention Encoder now only needs to learn an embedding-based representation of the task-specific delta target state $\Delta x_{t+1}^{*}$, which can be preprocessed by the pretrained PIDM Backbone. Meanwhile, the original output-layer (Prediction Decoder) is removed, and the concatenated outputs of the PIDM Backbone and Intention Encoder are passed in to a randomly initialized Action Synthesizer that synthesizes the final action $a_{t}$. PIDM is used in the critic via an almost identical architecture, with the only difference that the Action Synthesizer in the actor is replaced with a Value Synthesizer that outputs a scalar value estimation optimized with MSE loss.
@@
Intact RL setup: Except for the architectures of the actor and critic networks and the way the weights are initialized, our method does not require any modifications to either the POMDP (reward, curriculum design, observations, actions, and terminations) or to the PPO update rules, hyperparameters. The task-dependent Intention Encoder and Action Synthesizer can adapt to any configuration and dimension of the input and output. Therefore, the feasibility of handling arbitrary tasks is not limited. Every parameter in the pretrained PIDM remains trainable during the RL process. In this way, we allow task-specific dynamics to be learned during policy optimization, which eases the burden of attempting to exhaustively cover all possible transitions in the pretraining dataset.
@@ 5.1 Reinforcement Learning Tasks
We test our method on 9 RL environments with 3 distinct embodiments: a) 2 blind tasks (velocity-tracking locomotion (rudin2022learning) and pedipulation (arm_pedipulate_2024)) and 5 perceptive tasks (parkour walk, climb up, climb down, crouch, and jump (hoeller_anymal_2023)) with ANYmal-D (hutter_anymal_2016), b) velocity-tracking locomotion task with Unitree Go1 quadrupedal robot (default implementation in mittal2025isaac), and c) velocity-tracking locomotion task with Unitree G1 humanoid robot (default implementation in mittal2025isaac). All training is performed in Isaac Lab (mittal2025isaac).
[…]
@@ 5.2 Dynamics Knowledge Probing in Vanilla RL Policy Networks
(b) Errors of dynamics prediction based on representations from different layers, Pedipulation task.
Figure 4: Experiment to probe dynamics knowledge in vanilla policy networks. We analyze the correlation between intermediate representations between layers and the future joint state $q_{t+1}$. The zero-order extrapolation in (b) is a reference of the accuracy of always predicting $q_{t}=q_{t+1}$. Shaded areas indicate the standard deviation over 5 RL runs.
As a means of empirically studying our initial Hypothesis 1, we examine policy network model checkpoints from some task-specific RL process. For each checkpoint, we collect a number of observation-action pairs from the rollout distribution of the policy corresponding to that checkpoint. We then execute the mean action and record the resulting change in joint angles $\Delta q_{t+1}$. Meanwhile, we collect the intermediate representations from all three hidden layers, as illustrated in Figure 4(a).
We investigate how well the network understands at a specific layer what the consequences of its action will be by fitting a lightweight MLP to regress $\Delta q_{t+1}$ based on the tuple consisting of raw observation and intermediate representation corresponding to that layer. The lower final prediction error of a certain configuration indicate a better understanding.
Results of Pedipulation task is shown in Figure 4(b), and results of Locomotion task is shown in Figure 8. In the vertical direction, the correlation between the future state diminishes as we progress deeper into the network. This highlights the analogy of trained vanilla MLP policy networks to classic control from Section 4.1, where the model first forms an intent on the target state and subsequently computes the inverse dynamics to determine the required action. For more details on the experiment, we refer the readers to Appendix A.2.
@@ 5.3 Pretraining the Proprioceptive Inverse Dynamics Model
Figure 5: PIDM training and dataset analysis of ANYmal D: For the pretraining dataset we visualize its (a) coverage (green) compared to the initial exploration stages in RL (red and blue) using an UMAP projection and (b) the sample distribution of absolute action magnitudes $|a_{t}-q_{t}|$ over different terrains. Finally, in (c) we show the resulting PIDM accuracy across the entire action range as absolute joint errors $|\hat{q}_{t+1}-q_{t+1}|$ and also normalized by the action magnitude.
In this subsection, we describe how to obtain a pretrained PIDM model and analyze both the dataset distributions and the model’s accuracy, using ANYmal D as an example. We first analyze the quality of the data collected using the exploration-based strategy outlined in Section 4.3. In addition to the previously mentioned standard data augmentations (e.g., mass randomization, random noise, symmetry), we collect data on either or both flat and basic rough terrain generated with Perlin noise (miki_learning_2022; lee2020learning). In Figure 5(a) we plot samples of ANYmal D from the flat-terrain environments along with samples from the learning process of Pedipulation and Locomotion tasks, which are trained solely on flat terrain. Using UMAP (mcinnes2018umap), we project the proprioceptive observations $x$ for our collected dataset into 2D, and the observations from the first $100$ iterations of RL training for pedipulation and locomotion. We can thus validate that we obtain good coverage of, and beyond, the initial stages of the RL training process, which aligns with the goals outlined in Section 4.2.
[…]
This indicates the considerable difficulty in training a PIDM with high accuracy, which we attribute to the vast transition space, partial observability, and the lack of inductive bias in MLPs. For a detailed discussion, see Appendix A.4. While accurate data-driven modeling of robot locomotion may be achievable with models that are many orders of magnitude larger (xu2025neural) than those used here, actor and critic networks in motion-policy learning have traditionally been extremely lightweight. Moreover, large models are known to make reinforcement learning substantially more challenging (ota2021training; li2023survey). As a result, significantly increasing model capacity raises concerns about whether existing methods can still be applied without modification. For these reasons, we choose to keep the model size to millions of parameters which is much closer to that of the vanilla MLPs used in prior works. Although they may not seem too accurate for an inverse-dynamics solving, we will demonstrate in the following subsection that a pretrained module of such accuracy can already significantly enhance RL training. We further present a study of relation between error level of PIDM and RL performance in Appendix A.8 to verify that a trend of positive correlation between the accuracy of pretrained PIDM and resulted gain in RL performance can be observed.
@@ 5.4 Quantitative experiments
Metric Method ANYmal D Go1 G1 Avg.
Loco- Pedipu- Parkour Climb Climb Crouch Jump Loco- Loco-
motion lation Walk Up Down motion motion
Final perf. (%, $\uparrow$) Vanilla MLP +0.5 +0.2 -0.8 0.0 +6.5 +1.8 +11.1 +0.6 -0.2 +2.2
increase PIDM (Pretrained) +10.1 +6.3 +0.7 0.0 +27.7 +1.8 +5.9 +3.6 +10.0 +7.3
Num. iters. (%, $\downarrow$) Vanilla MLP -28.7 +5.0 -18.7 +22.5 -11.4 -29.5 -53.0 +2.2 -46.7 -17.6
to converge PIDM (Pretrained) -33.1 -42.0 -35.3 -20.6 -43.2 -57.3 -49.3 -17.7 -34.0 -36.9
@@
Figure 6: Evolution of the main performance metric during training for Locomotion, Pedipulation, Climb Up and Climb Down tasks with ANYmal D. The shaded areas denote standard deviations across five seeds.
In each experiment, we compare three methods: (i) the vanilla 4-layer MLP, (ii) our PIDM architecture with randomly initialized weights, and (iii) our PIDM architecture with pretrained weights. The utility of our method (i.e. using pretrained weights) is indicated by the comparison between (ii) and (iii). The performance of the vanilla 4-layer MLP is included only as a reference. Results are averaged over five runs with different random seeds, except for some individual cases mentioned in Appendix Table 8. To note is that we did not tune the learning parameters (learning rate, entropy coefficient, etc.) of the tasks, which were chosen for optimal performance of the original vanilla MLP. We merely used our architecture as a drop-in replacement. Therefore, it is possible that the performance of the proposed method can be further improved with additional tuning of these hyperparameters, which would have happened if the problem design had used our proposed architecture as a starting point.
@@
The selection of the main performance indicator varies across tasks (for details, see Appendix A.6). The results across all nine tasks are presented in Table 1. For some tasks, we also plot the evolution of the main performance indicator during training in Figure 10. The PIDM architecture with random weight initialization, PIDM (Random Init), generally lags behind the vanilla MLPs due to a larger model size and input dimension (inclusion of history). However, with the proposed pretraining strategy, PIDM (Pretrained) not only consistently outperforms PIDM (Random Init) in all metrics, but also significantly surpasses the performance of the vanilla MLP in 7 out of 9 tasks. Matching the performance of the MLP with our PIDM architecture is a secondary goal that could be potentially achieved by exhaustive tuning of the model and proprioceptive input. The key takeaway is the comparison between the randomly initialized and pretrained architectures. Compared with the vanilla MLP, PIDM (Pretrained) demonstrates an average improvement of 5.05.0% on final performance, and 18.818.8% on sample efficiency. When compared with PIDM (Random Init), the proposed PIDM (Pretrained) showcases an improvement of 7.37.3% on final performance, and enhances sample efficiency by a margin of 36.936.9%. We also note that despite the PIDM never having experienced the complex terrains used in the parkour task (see Figure 1), it quickly adapts to the new task-specific dynamics during RL training.
@@ 5.5 Ablations
Metric Method Anymal D
Climb Climb
@@
Metric Method Anymal D
Climb Climb
@@
We also perform 2 ablations with Climb Up and Climb Down tasks of ANYmal D, to motivate some of our design choices. First, in Table 2 we analyze the initialization strategy of choosing to pretrain either or both of the actor and critic networks. We note that only pretraining either the actor or critic generally still improves mean performance, but at the cost of increased instability, underscored by larger variation across runs (see Figure 12 in the Appendix). Second, we also ablate the data source used to pretrain the PIDM. We compare using exploration-based data versus samples from the initial stages of RL training a policy (in this case, pedipulation, as shown in Table 3. Notably, both datasets significantly outperform random initialization. Our approach provides an extra margin of improvement and theoretically adapts better to downstream tasks without overcomplicating the pipeline, as pretraining is performed only once.
@@ 5.6 Weight Update Magnitude
Figure 7: Network weight update magnitude comparison in the PIDM structured actor network during ANYmal D pedipulation RL training. In each submodule, the update of each linear layer weight is indicated by the mean absolute change per parameter, which is then averaged over all layers.
We compare the network weight update magnitudes between PIDM (Pretrained) and PIDM (Random Init) during the first 100 iterations of RL in Figure 7. We find that not only does the model exhibit smaller updates per iteration in the pretrained PIDM backbone, but this also results in smaller updates in the randomly initialized upstream Intention Encoder and downstream Action Synthesizer. This finding suggests that our pretrained weights lie closer to the desired local minimum and is an indicator that the optimization process can properly leverage this fact. For more examples, see Figure 15 in Appendix.
@@ 6 Conclusion
To summarize, we have presented a method for warm-starting the RL training process in actor-critic algorithms, targeted for robotic motion control. Our proposed approach leverages a network architecture based on a Proprioceptive Inverse Dynamics Model (PIDM) that is pretrained using exploration-based data from a specific robot embodiment. Our modular architecture functions as a drop-in replacement, without hyperparameter tuning, for any task on the pretrained robot embodiment. We demonstrate on 9 diverse RL environments with 2 quadrupedal robots and 1 humanoid robot that we can improve the final performance by $\mathbf{7.3}$%, and enhance sample efficiency by $\mathbf{36.9}$%. We also provide ablation studies to motivate our design choices and extensive empirical insights into the inner workings of our method. Future work can focus on optimizing model design to reduce the network size further and incorporating network architectures that are more adept at working with time-series data.